OpenStudy (anonymous):

Expeess the series 1+4+9+16 using sigma notation

OpenStudy (empty):

Ahhh this is a fun one! Any ideas? Give it your best shot and guess and I'll help you understand. I remember when I first saw sigma notation I didn't understand it either, but through guessing I can understand where you need help in understanding.

OpenStudy (skullpatrol):

want a hint?

OpenStudy (anonymous):

Yes

OpenStudy (empty):

Not even gonna try? :(

OpenStudy (skullpatrol):

$$\Huge1^2 = 1$$$$\Huge2^2 = 4$$$$\Huge3^2 = 9$$$$\Huge4^2 =16$$

OpenStudy (anonymous):

Do i add 1+4+9+16 by any chance? Sorry i dont remember how to do this

OpenStudy (skullpatrol):

yes

OpenStudy (skullpatrol):

Sigma notation is a shorthand way of writing: n + (n+1) + (n+2) + ...

OpenStudy (skullpatrol):

\[\sum_{1}^{4}\]

OpenStudy (skullpatrol):

"Sigma" means "sum"

OpenStudy (anonymous):

Oh ok.

OpenStudy (skullpatrol):